# Surrogate Models for Vibrational Entropy Based on a Spatial Decomposition

### 21 Feb 2024 v2

@@ Appendix B An alternative derivation of entropy
The DOS can be comprehended in the operational sense that if we integrate $\Omega(\omega)$ over a frequency range $\omega_{1}$ to $\omega_{2}$, we therefore obtain the total number of states within that frequency range [24]
@@
In the case where the defect does not add additional degrees of freedom to the lattice, e.g. for a substitutional defect, we have
@@ Appendix C Fast contour integration
Starting in the $s$-plane, we first map the annulus to a rectangle in the $t$-plane with vertices ±$\pm K$ and ±$\pm K+iK^{\prime}$, where $K,K^{\prime}$ denote the complete elliptic integrals using a logarithmic transformation
@@
where $K$ denotes the complete elliptic integral of the first kind. Next, the rectangle in the $t$-plane is mapped to the upper half-plane in the $u$-plane by the Jacobian elliptic function
@@ A Möbius transformation is then applied to map the upper half-plane to the $z$-plane
It is worthwhile mentioning that an additional computational cost for evaluating (C) stems from the computation of $\partial H(u)/\partial u_{n}$. In the locality test shown in Section 3, one requires to compute $\partial H(u)/\partial u_{n}$ with respect to all sites $n$, which results in computing the Jacobian of the Hessian matrix. We leverage the sparsity of the Hessian matrix to compute its Jacobian tensor. The sparsity pattern of a differential matrix in advance can greatly streamline computations. This is achieved by compressing the sparse matrix into a denser format, allowing for the processing of non-zero entries with significantly fewer function calls than previously necessary. This technique uses a strategy from graph theory [21, 29], aiming to combine columns with non-overlapping non-zero elements, referred to as structurally orthogonal columns into single groups, thus reducing the total number of groups needed.
@@ Appendix D Atomic Cluster Expansion
We provide an overview of the Atomic Cluster Expansion (ACE) potential and refer to [1, 12, 5] for more detailed construction and discussion. The ACE site potential can be systematically formulated as an atomic body-order expansion for a given correlation order $N\in\mathbb{N}$,
$V_{\text{ACE}}\big{(}\{\mathbf{y}_{j}\}\big{)}=\sum_{N=0}^{\mathscr{N}}\frac{1}{N!}\sum_{j_{1}\neq\dots\neq j_{N}}V_{N}(\mathbf{y}_{j_{1}},\ldots,\mathbf{y}_{j_{N}}),$ (D.1)
@@
where $\phi_{{\bm{n}}{\bm{\ell}}{\bm{m}}}(\mathbf{y}):=P_{n}(y)Y_{{\bm{\ell}}}^{{\bm{m}}}(\hat{\mathbf{y}}),$ with $\mathbf{y}\in\mathbb{R}^{d}$, $y=|\mathbf{y}|$, and $\hat{\mathbf{y}}=\frac{\mathbf{y}}{y}$. The functions $Y_{\ell}^{m}$ denote the complex spherical harmonics for $\ell=0,1,\ldots$, and $m=-\ell,\ldots,\ell$, and $P_{n}(r)$ represents the radial basis functions for $n=0,1,\ldots$. These basis functions are further symmetrized to ensure permutation invariance,
$\tilde{V}_{N}=\sum_{({\bm{n}},{\bm{\ell}},{\bm{m}})\text{ ordered}}c_{{\bm{n}}{\bm{\ell}}{\bm{m}}}\sum_{\sigma\in S_{N}}\phi_{{\bm{n}}{\bm{\ell}}{\bm{m}}}\circ\sigma,$
where $S_{N}$ signifies the set of all permutations, $({\bm{n}},{\bm{\ell}},{\bm{m}})$ denote the lexicographically arranged tuples. Thereafter, we need to incorporate invariance under rotations and point reflections,
$\mathcal{B}_{{\bm{n}}{\bm{\ell}}i}=\sum_{{\bm{m}}\in\mathcal{M}_{{\bm{\ell}}}}\mathcal{U}^{{\bm{n}}{\bm{\ell}}}_{{\bm{m}}i}\sum_{\sigma\in S_{N}}\phi_{{\bm{n}}{\bm{\ell}}{\bm{m}}}\circ\sigma\quad\text{with}\quad\mathcal{M}_{{\bm{\ell}}}=\{\mu\in\mathbb{Z}^{N}\mid-\ell_{\alpha}\leq\mu_{\alpha}\leq\ell_{\alpha}\},$ (D.2)
where the coefficients $\mathcal{U}^{{\bm{n}}{\bm{\ell}}}_{{\bm{m}}i}$ are given in [1, Lemma 2]. We usually aim to evaluate the sum over all N-neighbour clusters as follows:
$\mathcal{B}_{{\bm{n}}{\bm{\ell}}i}\big{(}\{\mathbf{y}_{j}\}\big{)}=\sum_{1\leq j_{1}\leq\cdots\leq j_{N}\leq J}\mathcal{B}_{{\bm{n}}{\bm{\ell}}i}(\mathbf{y}_{j_{1}},\cdots,\mathbf{y}_{j_{N}}),$ (D.3)
which scales as $\binom{J}{N}$ and was shown in [1] to be computationally inefficient. We can then use a “density trick” technique [2, 12, 39] to transform this basis into one that is computational efficient with cost that scales linearly with N. The alternative basis is
$B_{{\bm{n}}{\bm{\ell}}i}(\{\mathbf{y}_{j}\})=\sum_{m\in\mathcal{M}_{{\bm{\ell}}}}\mathcal{U}^{{\bm{n}}{\bm{\ell}}}_{{\bm{m}}i}A_{{\bm{n}}{\bm{\ell}}{\bm{m}}}(\{\mathbf{y}_{j}\})\quad\text{ with the correlations}\quad A_{{\bm{n}}{\bm{\ell}}{\bm{m}}}(\{\mathbf{y}_{j}\}):=\prod_{\alpha=1}^{N}\sum_{j=1}^{J}\phi_{{\bm{n}}_{\alpha}{\bm{\ell}}_{\alpha}{\bm{m}}_{\alpha}}(\mathbf{y}_{j}),$ (D.4)
which avoids the cost of summing over all order $N$ clusters inside an atomic neighbourhood as well as the $N!$ cost for symmetrising the basis.
We can now describe the resulting basis by
$\mathbf{B}_{N}:=\big{\{}B_{{\bm{n}}{\bm{\ell}}i}\mid({\bm{n}},{\bm{\ell}})\in\mathbb{N}^{2N}\text{ ordered},~{}\sum_{\alpha}{\bm{\ell}}_{\alpha}\text{ even},i=1,\ldots,n_{{\bm{n}}{\bm{\ell}}}\big{\}},$
where $n_{{\bm{n}}{\bm{\ell}}}$ is the number of basis functions for the chosen $({\bm{n}},{\bm{\ell}})$ [1]. After building the finite symmetric polynomial basis set $\mathbf{B}\subset\bigcup_{N=1}^{\mathscr{N}}\mathbf{B}_{N}$, every choice of parameters $\mathbf{c}=(c_{B})_{B\in\mathbf{B}}$ defines a site potential
$V^{\text{ACE}}(\{\mathbf{y}_{j}\};\mathbf{c})=\sum_{B\in\mathbf{B}}c_{B}B(\{\mathbf{y}_{j}\})$ (D.5)
Due to the completeness of this representation, as the approximation parameters (like body-order, cut-off radius, and expansion precision) approach infinity, the model can represent any arbitrary potential [1].

### 18 Mar 2024 v3

@@ Appendix B An alternative derivation of entropy
The DOS can be comprehended in the operational sense that if we integrate $\Omega(\omega)$ over a frequency range $\omega_{1}$ to $\omega_{2}$, we therefore obtain the total number of states within that frequency range [12],
@@
In the case where the defect does not add additional degrees of freedom to the lattice, e.g. for a substitutional impurity, we have
@@ Appendix C Fast contour integration
We first map the annulus to a rectangle with vertices ±$\pm K$ and ±$\pm K+iK^{\prime}$, using a logarithmic transformation
@@
where $K,K^{\prime}$ denote the complete elliptic integrals. Next, the rectangle is mapped to the upper half-plane in the $u$-plane by the Jacobian elliptic function
@@ A Möbius transformation is then applied to map the upper half-plane to the $z$-plane
It is worthwhile mentioning that an additional computational cost for evaluating (C) stems from the computation of $\partial H(u)/\partial u_{n}$. In the locality test shown in Section 3, one requires to compute $\partial H(u)/\partial u_{n}$ with respect to all sites $n$, which results in computing the Jacobian of the Hessian matrix. We leverage the sparsity of the Hessian matrix to compute its Jacobian tensor. The sparsity pattern of a differential matrix in advance can greatly streamline computations. This is achieved by compressing the sparse matrix into a denser format, allowing for the processing of non-zero entries with significantly fewer function calls than previously necessary. This technique uses a strategy from graph theory [26, 33, 47], aiming to combine columns with non-overlapping non-zero elements, referred to as structurally orthogonal columns into single groups, thus reducing the total number of groups needed.
@@ Appendix D Atomic Cluster Expansion
We provide a brief overview of the Atomic Cluster Expansion (ACE) potential and refer to [1, 15, 48, 50] for more detailed construction and discussion. The ACE site potential can be systematically formulated as an atomic body-order expansion for a given correlation order $N\in\mathbb{N}$,
$V_{\text{ACE}}\big{(}\{\mathbf{y}_{j}\}_{j=1}^{J}\big{)}=\sum_{N=0}^{\mathscr{N}}\frac{1}{N!}\sum_{j_{1},\dots,j_{N}=1}^{J}V_{N}(\mathbf{y}_{j_{1}},\ldots,\mathbf{y}_{j_{N}}),$ (D.1)
@@
[…]
This parameterization is already invariant under permutations (or, relabelling) of the input structure $\{{\bf y}_{j}\}_{j}$. One then symmetrizes the tensor product basis with respect to the group $O(3)$, employing the representation of that group in the spherical harmonics basis. This results in a linear parameterization
$V_{\text{ACE}}\big{(}\{\mathbf{y}_{j}\}_{j=1}^{J}\big{)}$ $=\sum_{{{\bm{n}}{\bm{\ell}}q}}c_{{{\bm{n}}{\bm{\ell}}q}}B_{{\bm{n}}{\bm{\ell}}q}(\{\mathbf{y}_{j}\}),$
$\text{where}\quad B_{{\bm{n}}{\bm{\ell}}q}(\{\mathbf{y}_{j}\})$ $=\sum_{{\bm{m}}\in\mathcal{M}_{{\bm{\ell}}}}\mathcal{C}^{{\bm{n}}{\bm{\ell}}q}_{{\bm{m}}}{\bm{A}}_{{\bm{n}}{\bm{\ell}}{\bm{m}}}(\{\mathbf{y}_{j}\}),$
$\text{and}\quad{\bm{A}}_{{\bm{n}}{\bm{\ell}}{\bm{m}}}(\{\mathbf{y}_{j}\})$ $=\prod_{\alpha=1}^{N}\sum_{j=1}^{J}\phi_{{\bm{n}}_{\alpha}{\bm{\ell}}_{\alpha}{\bm{m}}_{\alpha}}(\mathbf{y}_{j}).$
The $q$-index ranges from $1$ to $n_{\bm{n}\bm{\ell}}$ and enumerates the number of all possible invariant couplings through the generalized Clebsch–Gordan coefficients $\mathcal{C}^{{\bm{n}}{\bm{\ell}}q}_{{\bm{m}}}$. The This parameterization was originally devised in [17]. Its approximation properties and computational complexity are analyzed in detail in [1, 2, 6, 10]. The implementation we employ in our work is described in [50]. The latter reference also described the details of the choice of radial basis $P_{n}$.
To complete the description of our parameterization, we select a finite subset of the basis,
$\mathbf{B}:=\Big{\{}B_{{\bm{n}}{\bm{\ell}}q}\,\Big{|}\,$ $({\bm{n}},{\bm{\ell}})\in\mathbb{N}^{2N}\text{ ordered},~{}\sum_{\alpha}\ell_{\alpha}\text{ even},\sum_{\alpha}m_{\alpha}=0,$
$q=1,\ldots,n_{{\bm{n}}{\bm{\ell}}},\sum_{\alpha}\ell_{\alpha}+n_{\alpha}\leq D_{\rm tot},N\leq\mathscr{N}\Big{\}},$
where the two approximation parameters are the correlation order $\mathscr{N}$ and the total degree $D_{\rm tot}$. With this selection of the basis we can write our ACE parameterization more convenient as
$V^{\text{ACE}}(\{\mathbf{y}_{j}\};\mathbf{c})=\sum_{B\in\mathbf{B}}c_{B}B(\{\mathbf{y}_{j}\}),$ (D.2)
where ${\bf c}=(c_{B})_{B\in{\bf B}}$ are the model parameters. Due to the completeness of this representation, as the approximation parameters (like body-order, cut-off radius, and expansion precision) approach infinity, the model can represent any arbitrary potential [1].